# Rethinking Metrics and Benchmarks of Video Anomaly Detection

### v1

@@ Abstract
Video Anomaly Detection (VAD), which aims to detect anomalies that deviate from expectation, has attracted increasing attention in recent years. Existing advancements in VAD primarily focus on model architectures and training strategies, while devoting insufficient attention to evaluation metrics and benchmarks. In this paper, we rethink VAD evaluation protocols through comprehensive experimental analyses, revealing three critical limitations in current practices: 1) existing metrics are significantly influenced by single annotation bias; 2) current metrics fail to reward early detection of anomalies; 3) available benchmarks lack the capability to evaluate scene overfitting. To address these limitations, we propose three novel evaluation methods: first, we establish averaged AUC/AP metrics over multi-round annotations to mitigate single annotation bias; second, we develop a Latency-aware Average Precision (LaAP) metric that rewards early and accurate anomaly detection; and finally, we introduce two hard normal benchmarks (UCF-HN, MSAD-HN) with videos specifically designed to evaluate scene overfitting. We report performance comparisons of ten state-of-the-art VAD approaches using our proposed evaluation methods, providing novel perspectives for future VAD model development. Anonymous links for data and code are provided.
1 Introduction
@@
Video Anomaly Detection (VAD) aims to identify abnormal frames deviating from expectations in video sequences, playing a pivotal role in applications such as intelligent surveillance, traffic management, and healthcare [1, 2]. In recent years, this task has garnered increasing attention, with existing methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12] achieving satisfactory performance on several large-scale datasets [13, 14, 15] through weakly-supervised paradigms. Similar to other domains in machine learning, the advancement of VAD correlates with evaluation metrics and datasets. Despite its importance, this direction remains understudied. Inspired by similar studies in other domains [16, 17], we conduct a comprehensive analysis of the methods of evaluating VAD algorithms. Our investigation reveals that existing metrics and benchmarks fail to reliably and comprehensively evaluate VAD algorithms. We summarize the limitations in terms of annotation, metric, and benchmarking data as follows.
@@
For annotation, existing evaluation methods are generally limited to leveraging single-round annotations, where each video is annotated only once. However, as illustrated in Figure 1(a), individual annotations may exhibit bias - one tends to favor the most salient abnormal behavior frames while the other may favor more comprehensive event segments. The discrepancy of annotations could compromise the reliability of metric computations. [18] preliminarily reveals this phenomenon in the UCF-Crime dataset [13] by providing an alternative annotation. Some works [19, 20] also mention the label noise in UCF-Crime. However, these works lack systematic dataset analysis and focus exclusively on one dataset.
@@
For metric design, the most commonly used metrics in VAD are AUC (Area Under receiver operating characteristic Curve) and AP (Average Precision). These metrics treat VAD as a frame-wise binary classification task, which is insensitive to the temporal position of predictions. In safety-critical applications such as emergency response or fall detection systems, low anomaly detection latency  [^1] The latency in this paper denotes how rapidly scores rise after an anomaly occurs rather than computational throughput. is vital for triggering actionable alerts. Figure 1(b) illustrates two detection curves achieving identical AUC/AP, but the first curve exhibits an earlier detection compared to the second. Although there are metrics that address detection latency in time series anomaly detection [21, 22] and continual VAD [23], they assume anomalies as isolated points rather than temporal events, rendering them inapplicable to generic VAD evaluation.
@@
For benchmarking data, current large-scale datasets primarily collect videos from the internet or movies with various scenes. As exemplified in Figure 1(c), a typical limitation arises when a trained model is tested on scenes that are different from the training data. These test data are relatively easy for the model to detect and cannot tell whether the model is overfitted to the scene (e.g., taking a cramped grocery as an anomaly). The ideal test data should comprise videos of the same scene with different events (e.g., shopping in a cramped store peacefully). Though some datasets [24, 25, 26, 27, 28] collected by manual recording or 3D animations contain different behaviors in the same scene, they either lack severe anomalies (e.g., explosions), exhibit domain gaps from real-world data, or suffer from the limited scale of dataset samples.
@@
In this paper, we focus on better evaluation methods for VAD algorithms. First, we systematically analyze the three aforementioned limitations on three large-scale VAD datasets (UCF-Crime [13], XD-Violence [15], MSAD [14]) through multi-round re-annotation (i.e., each round is labeled by different annotators) and exploratory experiments. Second, we propose averaged AUC/AP metrics to mitigate annotation bias and a novel Latency-aware Average Precision (LaAP) metric to evaluate the latency of VAD algorithms. The averaged AUC/AP calculate the AUC/AP of each round of annotation and take the average. The LaAP replaces AP’s Precision-Recall curve with a Precision-LaRecall curve, where LaRecall rewards earlier detections of anomalies through a time-decaying weighting mechanism. This metric also maintains the capability to evaluate model performance across varying thresholds. Third, we construct two synthetic benchmarks (UCF-HN & MSAD-HN) via image-to-video diffusion models [29, 30] to evaluate overfitting in VAD algorithms. UCF-HN and MSAD-HN comprise normal videos generated from keyframes of abnormal scenes in UCF-Crime and MSAD training sets, where high false alarm rates (FAR) on these data reflect overfitting.
@@
    We propose the annotation-averaged AUC/AP metrics and the LaAP metric, which mitigate annotation bias and reward early anomaly detection.
@@
2 Related Works
2.1 Video Anomaly Detection Metrics
@@
The majority of existing works employ AUC and AP as evaluation metrics. Here, AUC represents the area under the TPR (True Positive Rate)-FPR (False Positive Rate) curve, while AP denotes the area under the Precision-Recall curve. Compared to Accuracy and F1-score, both metrics comprehensively evaluate the performance under varying decision thresholds, with AUC being insensitive to class imbalance [34] and AP demonstrating better capability for positive instances [15]. Despite criticisms of these metrics [35, 34], they remain widely adopted in medical diagnosis [36], object detection [37], and recommendation systems [38]. In video anomaly detection, anomalies typically manifest as continuous multi-frame events, and early detection of such anomalies is often prioritized in practical applications. However, both AUC and AP treat predictions as independent instances, thereby neglecting the event-based nature of anomalies and exhibiting insensitivity to detection latency. While the NAB metric [22] in time-series anomaly detection and the APD metric [23] in continual learning VAD incorporate prediction latency, they model anomalies as discrete points and measure delays within a fixed time window, which overlook variations in duration of abnormal events, rendering them suboptimal for VAD evaluation. There are also metrics concerning spatiotemporal localization (e.g., TBDC [39], STAUC [40]) or multi-class classification (e.g., mAP@IoU [10]), but we focus on temporal detection performance in VAD. Moreover, existing metrics only utilize single-round annotation per test video, potentially introducing annotator bias as discussed in [18, 41, 42].
@@
Inspired by them, we propose the Latency-aware Average Precision (LaAP), which rewards early anomaly detection through an adaptive time window and nonlinear score functions, offering a novel perspective for VAD evaluation. We also conduct independent re-annotations for three large-scale datasets, where the averaging metric helps mitigate individual annotator bias.
2.2 Video Anomaly Detection Benchmarks
@@
Table 1: Comparisons of VAD datasets. Intra-Scene Pos&Neg denotes the inclusion of both normal and abnormal videos of the same scene.
Dataset Source # videos
Intra-Scene
Pos&Neg
@@
Current VAD benchmarks predominantly adopt in-domain evaluation, with few works [41, 42, 44, 45, 46] employing cross-domain approaches. The in-domain protocol evaluates models on test sets corresponding to their training data, whereas the cross-domain protocol evaluates generalization capability using test sets from different datasets. Table 1 compares existing VAD datasets. [25, 43, 24, 28, 26] manually record videos with actors performing abnormal events, but these datasets suffer from limited scale and lack severe anomaly events (e.g., explosions). [27] utilizes synthetic 3D animations to generate videos, but it introduces significant discrepancies from real-world data. [13, 15, 14] collect videos from the web and movies, greatly expanding anomaly categories and scale. However, unlike manually recorded datasets, they fail to provide both normal and abnormal videos of the same scene, making it challenging to determine whether the model has overfitted to scenes rather than the true anomaly patterns. For instance, a dataset might contain an abnormal video showing a van exploding at a gas station, but lack normal videos where the same van refuels and departs safely at the same gas station. An overfitting model might consider gas stations or vans as anomalies, yet both in-domain and cross-domain benchmarks cannot evaluate that.
To address these issues, we construct UCF-HN and MSAD-HN, generated via image-to-video diffusion models [29, 30], which synthesize normal behaviors under identical scenes to evaluate overfitting.
3 Analysis of Current Metrics and Benchmarks
3.1 Bias Prevails in Annotations
@@
Table 2: The overall Fleiss’ Kappa and the median of standard deviations for start time, duration, and end time (in seconds).
@@
The validity of evaluation metrics depends on the precision of annotations. In this section, we systematically analyze annotation discrepancies by re-annotating three large-scale datasets. Specifically, we select UCF-Crime [13], XD-Violence [15], and MSAD [14] datasets, among which MSAD represents a newer dataset featuring broader anomaly categories and higher video resolutions. Dataset details are provided in the supplementary Section C.
We re-annotate frame-level temporal boundaries for all abnormal videos in the test sets of these three datasets, strictly adhering to the original papers’ criteria. Each annotation round is assigned to a different annotator for independent labeling. This process yields four independent annotation rounds per dataset, including their vanilla annotations. Notably, for UCF-Crime, one round is from [18]. On UCF-Crime, we find that some abnormal videos have category ambiguity. Consequently, we perform additional categorical annotation on all abnormal videos of UCF-Crime.
We quantify the overall frame-level annotation consistency across three re-annotated datasets using Cohen’s Kappa [47] and Fleiss’ Kappa [48] in Figure 2 and Table 2. Both metrics range in $[-1,1]$ where values approaching 1 indicate perfect agreement. The results reveal that the lowest inter-annotator Cohen’s Kappa is 0.382, with overall Fleiss’ Kappa ranging from 0.51 to 0.68, indicating substantial annotation discrepancies. To identify the source of discrepancies, we analyze the distribution of standard deviations of start time, duration, and end time in Figure 2 and Table 2. Note that the start and end times correspond to the earliest and latest annotated abnormal frames in a video, and the duration refers to the total time of all annotated abnormal segments in a video. The results reveal that the deviations for the duration and end time are significantly higher than those for the start time, which indicates that annotators diverge in granularity (i.e., whether to label the most salient segments or more comprehensive events) and determination of ending (i.e., whether to label the aftermath).
Figure 3 visualizes the categorical discrepancies between the original annotation and ours in UCF-Crime. Videos originally labeled as Stealing, Shoplifting, and Burglary exhibit high rates of “misclassification” as Normal ($\geq 30\%$). We attribute this to inherent dependence on contextual information beyond the video. For instance, it is challenging to distinguish between the staff member’s routine actions and a shoplifter’s behavior of taking away a t-shirt in surveillance footage without knowing who the staff member is. Such ambiguous instances introduce confounding signals during training. Visualization examples of annotation bias are given in Section C.3.
@@
Figure 4: Heatmap of predicted anomaly positions of different models [3, 4, 5, 9, 10, 7, 8, 31, 32, 33] in abnormal intervals. Brighter areas indicate higher density.
3.2 Existing Methods Differ in Latency
In practical applications, accurate and early anomaly detection is preferable under equivalent model capability. However, current metrics treat VAD task as a frame-independent binary classification task while neglecting positional information in predictions.
Figure 4 illustrates the distribution of prediction latency across multiple models. We select prediction scores within abnormal intervals, binarize them, obtain the positions of positive instances, and normalize them using the length of the abnormal interval. The results reveal that certain models (e.g., RTFM [3], CLIP-TSA [8]) exhibit relatively uniform predictions, while some (PEL [9], Glance-VAD [32]) achieve higher positive concentrations in earlier stages, indicating differences exist in models’ capability of predicting latency.
3.3 Current Benchmarks Struggle to Evaluate Scene Overfitting
Table 3: Re-splitting strategy and performance of on the re-split ShanghaiTech.

### v2

@@ Abstract
Video Anomaly Detection (VAD), which aims to detect anomalies that deviate from expectation, has attracted increasing attention in recent years. Existing advancements in VAD primarily focus on model architectures and training strategies, while devoting insufficient attention to evaluation metrics and benchmarks. In this paper, we rethink VAD evaluation methods through comprehensive analyses, revealing three critical limitations in current practices: 1) existing metrics are significantly influenced by single annotation bias; 2) current metrics fail to reward early detection of anomalies; 3) available benchmarks lack the capability to evaluate scene overfitting of fully/weakly-supervised algorithms. To address these limitations, we propose three novel evaluation methods: first, we establish probabilistic AUC/AP (Prob-AUC/AP) metrics utlizing multi-round annotations to mitigate single annotation bias; second, we develop a Latency-aware Average Precision (LaAP) metric that rewards early and accurate anomaly detection; and finally, we introduce two hard normal benchmarks (UCF-HN, MSAD-HN) with videos specifically designed to evaluate scene overfitting. We report performance comparisons of ten state-of-the-art VAD approaches using our proposed evaluation methods, providing novel perspectives for future VAD model development. We release our data and code in https://github.com/Kamino666/RethinkingVAD.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
Video Anomaly Detection (VAD) aims to identify abnormal frames deviating from expectations in video sequences, playing a pivotal role in applications such as intelligent surveillance, traffic management, and healthcare [1, 2]. In recent years, this task has garnered increasing attention, with existing methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] achieving satisfactory performance on several large-scale datasets [16, 17, 18]. Similar to other domains in machine learning, the advancement of VAD correlates with evaluation metrics and datasets. Despite its importance, this direction remains understudied. Inspired by similar studies in other domains [19, 20], we conduct a comprehensive analysis of the methods of evaluating VAD algorithms. The scope of this study is delineated as the temporal localization of anomalies in videos under general and unconstrained scenes. Our investigation reveals that existing metrics and benchmarks fail to reliably and comprehensively evaluate VAD algorithms. We summarize the limitations in terms of annotation, metric, and benchmarking data as follows.
@@
For annotation, existing evaluation methods are generally limited to leveraging single-round annotations, where each video has only single hard label. However, individual annotations may exhibit bias. In Figure 1(a), one tends to favor the most salient abnormal behavior frames while the other may favor more comprehensive event segments. The discrepancy of annotations could compromise the reliability of metric computations. [21] preliminarily reveals this phenomenon in the UCF-Crime dataset [16] by providing an alternative annotation. Some works [22, 23] also mention the label noise in UCF-Crime. However, these works lack systematic dataset analysis and focus exclusively on one dataset.
@@
For metric design, the most commonly used metrics in this field are AUC (Area Under receiver operating characteristic Curve) and AP (Average Precision). These metrics treat VAD as a frame-wise binary classification task, which is insensitive to the temporal position of predictions. In safety-critical applications such as emergency response or fall detection systems, low anomaly detection latency  [^1] The latency in this paper denotes how rapidly scores rise after an anomaly occurs rather than computational throughput. is vital for triggering actionable alerts. Figure 1(b) illustrates two detection curves achieving identical AUC/AP, but the first curve exhibits an earlier detection compared to the second. Although there are metrics that address detection latency in time series anomaly detection [24, 25] and continual VAD [26], they assume anomalies as isolated points rather than temporal events, rendering them inapplicable to generic VAD evaluation.
@@
For benchmarking data, current large-scale datasets primarily collect videos from the internet or movies with various scenes. As exemplified in Figure 1(c), a common issue arises when the test scenes are visually similar to those in the training data but contain different events. Such test data are too easy for the model and fail to reveal whether it has overfitted to the scene (e.g., treating a crowded grocery store as abnormal). The ideal test data should comprise videos of the same scene with different events (e.g., shopping in a cramped store peacefully). However, these data are not available on the web. Though some datasets [27, 28, 29, 30, 31] collected by manual recording or 3D animations contain different behaviors in the same scene, they either lack severe anomalies (e.g., explosions), exhibit domain gaps from real-world data, or suffer from the limited scale of dataset samples.
@@
In this paper, we focus on better evaluation methods for VAD algorithms. First, we systematically analyze the three aforementioned limitations on three large-scale VAD datasets (UCF-Crime [16], XD-Violence [18], MSAD [17]) through multi-round re-annotation (i.e., each round is labeled by different annotators) and exploratory experiments. Second, we introduce probabilistic AUC/AP metrics (Prob AUC/AP) to mitigate annotation bias and propose a novel Latency-aware Average Precision (LaAP) metric to evaluate the latency of VAD algorithms. The probabilistic AUC/AP aggregates multiple rounds of annotations into probability labels, then extends the confusion matrix to the cumulative probabilities, and calculate AUC/AP based on that. The LaAP replaces AP’s Precision-Recall curve with a Precision-LaRecall curve, where LaRecall rewards earlier detections of anomalies through a time-decaying weighting mechanism. Third, we construct two synthetic benchmarks (UCF-HN & MSAD-HN) via image-to-video diffusion models [32, 33] to evaluate overfitting in VAD algorithms. UCF-HN and MSAD-HN comprise normal videos generated from keyframes of abnormal scenes in UCF-Crime and MSAD training sets, where high false alarm rates (FAR) on these data reflect overfitting.
@@
    We introduce the Prob-AUC/AP metrics and propose the LaAP metric, which mitigate annotation bias and reward early anomaly detection.
@@
II Related Works
II-A Video Anomaly Detection Metrics
@@
The majority of existing works employ AUC and AP as evaluation metrics. Here, AUC represents the area under the TPR (True Positive Rate)-FPR (False Positive Rate) curve, while AP denotes the area under the Precision-Recall curve. Both metrics comprehensively evaluate the performance under varying decision thresholds, with AUC being insensitive to class imbalance [37] and AP demonstrating better capability for positive instances [18]. In video anomaly detection, anomalies typically manifest as continuous multi-frame events, and early detection of such anomalies is often prioritized in practical applications. However, both AUC and AP treat predictions as independent instances, thereby neglecting the event-based nature of anomalies and exhibiting insensitivity to detection latency. While the NAB metric [25] in time-series anomaly detection and the APD metric [26] in continual learning VAD incorporate prediction latency, they model anomalies as discrete points and measure delays within a fixed time window, which overlook variations in duration of abnormal events, rendering them inapplicable for general VAD evaluation. Moreover, existing metrics only utilize single-round annotation, which would introduce annotator bias, as mentioned in some works [21, 38, 39] and systematically analyzed in this paper.
@@
To address these issues, we propose LaAP, which rewards early anomaly detection through an adaptive time window and nonlinear score functions, offering a novel perspective for VAD evaluation. We also conduct independent re-annotations for three large-scale datasets and introduce ProbAUC and ProbAP to mitigate individual annotator bias.
II-B Video Anomaly Detection Benchmarks
Table I compares existing VAD datasets. Some datasets [28, 40, 27, 31, 29] manually record videos with actors performing abnormal events, but they are limited in scale and lack severe anomaly events (e.g., explosions). In contrast, [30] obtains videos using synthetic 3D animations, which is able to provide fine-grained pixel-level labels. Nevertheless, it introduces significant discrepancies from real-world data, limiting its applicability.
To address these limitations, datasets like [16, 18, 17] collect videos from the web and movies, greatly expanding the range of anomaly categories and increasing dataset scale. UCF-Crime [16] contains 1,900 surveillance videos from the web, covering 13 crime-related categories. XD-Violence [18] is another large-scale dataset with 4,754 videos with an audio channel, which spans 6 violence-related categories. Its videos primarily come from movies, news, and user-uploaded content on the web, often involving significant editing. MSAD [17] is a more recent dataset that covers 14 distinct scenarios and includes 11 base categories, with videos of higher resolution. The anomalies in MSAD range from human-related events, such as people falling, to non-human-related anomalies. However, these datasets do not provide both normal and abnormal videos for the same scene, which makes it difficult to evaluate whether a model has overfitted to specific scenes rather than learning to detect true anomaly patterns. For example, a dataset might contain a video of a van exploding at a gas station but lacks a normal video showing the same van refueling and departing safely from the same location. An overfitting model might mistakenly treat the gas station or the van as anomalies.
In this work, we construct UCF-HN and MSAD-HN, generated via image-to-video diffusion models, which synthesize normal behaviors under identical scenes to evaluate scene overfitting.
II-C Synthesizing Videos in Video Anomaly Detection
Due to the sparsity of anomalies in reality, some works have explored data synthesis. GTACrash [41] acquires crash data from gaming environments, exhibiting limited physical realism. UBNormal [30] mploys 3D animation, demonstrating suboptimal human motion naturalness and environmental integration. OVVAD [42] employs a text-to-image diffusion model coupled with an image animation model for frame generation, resulting in conspicuous visual artifacts and temporal inconsistencies. Astrid et al. [43] and Narwade et al. [44] generate anomalies through cut-and-paste operations, leading to significant background inconsistencies. Apart from the issues related to the quality of synthesizing, these methods only focus on using the synthesized data for training.
In contrast, this work emphasizes using synthetic data to evaluate scene overfitting in scenarios where obtaining real-world data is extremely challenging. We also employ state-of-the-art video generation models and incorporate manual efforts to ensure high-quality synthesis.
III Analysis of Current Metrics and Benchmarks
III-A Bias Prevails in Annotations
@@
TABLE I: Comparisons of VAD datasets. Intra-Scene Pos&Neg denotes the inclusion of both normal and abnormal videos of the same scene.
Dataset Source # videos Intra-Scene Pos&Neg
@@
The validity of evaluation metrics depends on the precision of annotations. In this section, we systematically analyze annotation discrepancies by re-annotating three large-scale datasets: UCF-Crime, XD-Violence, and MSAD. We conduct frame-level re-annotation of temporal boundaries for all abnormal videos in the test sets, requiring annotators to strictly follow the official definitions of each dataset. For example, in UCF-Crime, the category “Burglary” is defined as: This event contains videos that show people (thieves) entering into a building or house with the intention to commit theft. It does not include use of force against people. Each annotation round is assigned to a different annotator for independent labeling. In total, we obtain four independent rounds of annotations per dataset, including the original ones. Notably, for UCF-Crime, one annotation round is adopted from [21].
@@
TABLE II: The overall Fleiss’ Kappa and the median of standard deviations for start time, duration, and end time (in seconds).
@@
During this process, we observe that some abnormal videos in UCF-Crime exhibit category ambiguity. To address this, we further perform categorical re-annotation of all abnormal videos in UCF-Crime. Importantly, annotators were not informed that only abnormal videos were selected, and the annotators responsible for categorical re-annotation were different from those performing temporal boundary labeling, thereby mitigating potential bias.
We quantify the overall frame-level annotation consistency using Cohen’s Kappa [45] and Fleiss’ Kappa [46] in Figure 2 and Table II. Both metrics range in [1,1][-1,1] where values approaching 1 indicate perfect agreement. The results reveal that the lowest inter-annotator Cohen’s Kappa is 0.382, with overall Fleiss’ Kappa ranging from 0.51 to 0.68, indicating substantial annotation discrepancies. To identify the source of discrepancies, we analyze the distribution of standard deviations of start time, duration, and end time in Figure 2 and Table II. Note that the start and end times correspond to the earliest and latest annotated abnormal frames in a video, and the duration refers to the total time of all annotated abnormal segments in a video.
Finding 1. The results reveal that the deviations for the duration and end time are significantly higher than those for the start time, which indicates that annotators diverge in granularity (i.e., whether to label the most salient segments or more comprehensive events) and determination of ending (i.e., whether to label the aftermath).
@@
For categorical ambiguity, we visualizes the discrepancies between the original annotation and ours in UCF-Crime in Fig. 3. Videos originally labeled as Stealing, Shoplifting, and Burglary exhibit high rates of “misclassification” as Normal ($\geq 30\%$). We attribute this to inherent dependence on contextual information beyond the video. For instance, it is challenging to distinguish between the staff member’s routine actions and a shoplifter’s behavior of taking away a t-shirt in surveillance footage without knowing who the staff member is. Such ambiguous instances would introduce confounding signals during modeling and evaluating. Visualization examples are given in the supplementary.
Finding 2. Categories whose recognition inherently relies on contextual knowledge beyond visual evidence are less suitable for evaluation for vision-based methods.
Therefore, in the following experiments on UCF-Crime, we exclude these three ambiguous categories.
III-B Existing Methods Differ in Latency
In practical applications, accurate and early anomaly detection is preferable under equivalent model capability. However, current metrics treat VAD task as a frame-independent binary classification task while neglecting positional information in predictions. Figure 4 illustrates the distribution of prediction latency across multiple models. We select prediction scores within abnormal intervals, binarize them, obtain the positions of positive instances, and normalize them using the length of the abnormal interval. The results reveal that certain models (e.g., RTFM [3], CLIP-TSA [8]) exhibit relatively uniform predictions, while some (PEL [9], Glance-VAD [35]) achieve higher positive concentrations in earlier stages.
Finding 3. Existing models exhibit different behaviors in prediction latency.
III-C Current Benchmarks Struggle to Evaluate Scene Overfitting
Existing VAD models trained on large-scale web-sourced datasets [16, 18, 17] with either weak or full supervision are prone to overfitting to training scenes. However, current benchmarks are insufficient to evaluate such overfitting, as their test sets mainly consist of scenes that differ substantially from those in the training data. This makes them relatively easy for the models, while leaving open the question of whether performance would remain robust on test scenes that closely resemble the training ones. For example, a production scenario may involve a crowded grocery store with a checkout counter on the right, which is included in the training data, whereas the existing test data only contains videos from mobile phone stores and clothing retailers.
Figure 4: Heatmap of predicted anomaly positions of different models in abnormal intervals. Brighter areas indicate higher density. The introducion of these models are in Sec. V-A.
TABLE III: Re-splitting strategy and performance on the re-split ShanghaiTech.